Displaying report 1-1 of 1.
Reports until 12:21, Friday 07 June 2024
GQuEST General
Torrey Cullen - posted 12:21, Friday 07 June 2024 - last comment - 13:38, Friday 07 June 2024(11638)
More filter cavity science

[Daniel,Torrey]

-1550 REFL path is not aligned because the REFL beam is missing the mirror high. This is not ideal but i've added a quarter inch spacer to highten the first REFL mirror and then tilt the beam back down to 4 in height. We are operating in an increased power mode as well, so need to adjust the ND filters in front of the 1811. We have ~50 mW on the refl path and the 1811 saturates at 55 uW. Replacing the 2 + .4 ND filters with a 3. Now have 1550 PDH sidebands.

-We have a 11 Hz sinewave on the error signals. This seems to be caused from cross talk on the EOM drives. We replaced the BNC cables that are providing the EOM signal with the homemade ultra shield BNC cable. This did not make a difference. 

-We then wrapped the seperate inductors making up the tank circuits on the EOM in multiple layers of aluminum foil in case the RF was eminating from there. This seems to have made a difference but I now suspect I ripped the solder connection on one of the circuits as there are no PDH sidebands on the 1550 path anymore.

Comments related to this report
Torrey Cullen - 13:38, Friday 07 June 2024 (11639)

The foil was shorting the tank circuit. Fixed.

Displaying report 1-1 of 1.